# Massive $\nu$s through the CNN lens: interpreting the field-level neutrino mass information in weak lensing

### v2

@@ Abstract
Modern cosmological surveys probe the Universe deep into the nonlinear regime, where massive neutrinos suppress cosmic structure. Traditional cosmological analyses, which use the 2-point correlation function to extract information, are no longer optimal in the nonlinear regime, and there is thus much interest in extracting beyond-2-point information to improve constraints on neutrino mass. Quantifying and interpreting the beyond-2-point information is thus a pressing task. We study the field-level information in weak lensing convergence maps using convolution neural networks. We find that the network performance increases as higher source redshifts and smaller scales are considered – investigating up to a source redshift of 2.5 and $\ell_{\rm max}\simeq 10^{4}$ – verifying that massive neutrinos leave a distinct effect on weak lensing. However, the performance of the network significantly drops after scaling out the 2-point information from the maps, implying that most of the field-level information can be found in the 2-point correlation function alone. We quantify these findings in terms of the likelihood ratio and also use Integrated Gradient saliency maps to interpret which parts of the map the network is learning the most from, finding that the network extracts information from the most overdense and underdense regions.
@@ 1 Introduction
One particularly promising method to measure neutrino mass is by studying weak gravitational lensing. This involves analyzing the distortions of light emitted from distant galaxies caused by the gravitational influence of LSS between the galaxy and the telescope. Traditionally, information is extracted from weak-lensing convergence maps using the power spectrum, or 2-point correlation function, which is optimal in the linear regime. However, upcoming surveys will probe progressively smaller, nonlinear, scales, in which regime the 2-point statistics are no longer guaranteed to be optimal. Various simulations have been developed to understand the effects of neutrinos on nonlinear scales [16, 17, 18, 19, 20, 21, 22], and have suggested that there is much information to be found on these nonlinear scales using higher-order statistics such as voids and wavelets. For example, when considering the 3-dimensional matter field it has been shown that higher-order statistics can improve neutrino mass constraints by a factor of $\sim 100$ compared to the power spectrum alone [23, 24, 25, 26, 27, 28, 29, 30], while for weak lensing more modest numbers are reported [31, 32, 33, 34, 35, 36]. Understanding and interpreting the source of this information is thus a pressing task.
@@ 2.1 Data
    Addition of Gaussian Noise: Gaussian noise with zero mean was added to mimic observational noise. We considered two standard deviations of $\sigma_{\kappa}=0.01$ and 0.1 (in the same units as $\kappa$), which is around the order of magnitude of noise expected for an LSST-like survey with a galaxy number density of $n_{\rm gal}\simeq 5\,{\rm arcmin}^{-2}$ and a shape noise of $\sigma_{\lambda}\simeq 0.3$ [32]. We also consider the noiseless maps.
@@
    Cutting in Fourier Space: The images were transformed to Fourier space and cut at various scales in factors of 2 of the Nyquist, $k_{\rm max}=2\pi,\pi,\pi/2,\pi/4,\pi/8,\pi/16\,({h/{\rm Mpc})^{2}}$, and then transformed back into configuration space. As the maps are produced using the flat-sky approximation, we use a 2-dimensional wavenumber, denoted $k$, to represent the scale. To connect with curved-sky intuition, we show the relation between $k$ and multipole $\ell$ in Figure 2 for the various source redshifts considered.
@@ 2.2 Neural Network
We modified the ResNet-18 for our applications to take only 1 channel of input, instead of the default 3. We also replaced the final layer of the network by 3 fully connected layers which output a scalar, followed by a dropout, LeakyReLU, and sigmoid layer. We then apply a binary cross entropy loss function to train a classifier between the 0eV and 0.1eV neutrino mass simulations. For optimization we used the ADAM optimizer with an exponential learning rate scheduler, using early stopping to avoid overfitting. We report results for a batch size of 32, but also found that using a batch size of 64 or 128 made no improvement to the model. We use the pre-trained ResNet-18 weights as the starting point for training, but then refit all the weights during training. The data set was divided into training, validation, and testing sets according to a 70:15:15 split.
@@ 3 Results
Figure 3: Top: classification accuracy evaluated on the test set as a function of $k_{\text{max}}$. Bottom: difference in log likelihood between a $M_{\nu}=0.1$ and $0\,{\rm eV}$ for observations with $M_{\nu}=0.1\,{\rm eV}$ (taken from the test set) as a function of $k_{\text{max}}$. Each panel and base color corresponds to a different source redshift $z$. The difference in color shades corresponds to the noise $\sigma_{\kappa}$, with lighter shades corresponding to higher noise. Solid lines represent the accuracy for a model trained on the full convergence maps, while dashed lines are for maps that have had the 2-point information scaled out. It can be seen for full maps that the accuracy and difference in likelihood are typically higher at smaller scales, higher redshifts, and lower noise. However, when the 2-point information is removed the accuracy and difference in likelihood are consistently significantly worse – almost $50\%$ and 0 respectively – indicating that much of the field-level information is contained in the 2-point alone.
@@
Figure 4: Visualization of the logarithm of the absoulte value of the Integrated Gradient saliency maps for the case of $z=2.5$, $k_{\text{max}}={\pi}/{2}$, and $\sigma_{\kappa}=0.01$ (without scaling out the 2-point information). Each row displays a different (mean-subtracted and variance-scaled) weak lensing convergence map in the left column, followed by three saliency maps generated by the best three models from the optuna trials. Higher numbers in the saliency map implies that network is paying more attention to that region of the map. The saliency maps show a wide variability in terms of regions of the map they focus on, in some cases focusing on underdense regions, and in other cases focusing on overdense regions.
@@
Figure 5: The pixel-averaged value of the logarithm of the absolute value of the Integrated Gradient saliency maps as a function of the (mean-subtracted and variance-scaled) convergence map pixel value $\tilde{\kappa}$ for the case of $z=2.5$, $k_{\text{max}}={\pi}/{2}$, and $\sigma_{\kappa}=0.01$ (without scaling out the 2-point information). Higher saliency values imply that the network is paying more attention to that region of the map, thus the network is paying most attention to extreme values of $\kappa$, i.e. the most overdense and underdense regions.
@@
The upper row of Figure 3 shows the classification accuracy evaluated on the test set for the different scale cuts $k_{\rm max}$, source redshifts $z$, and noise levels $\sigma_{\kappa}$, as well as the effect of scaling out the 2-point information. It can be seen that the accuracy of the network improves as smaller, nonlinear, scales are included; this is to be expected as the effect of massive neutrinos becomes more prominent on small scales. It can also be seen that the accuracy is higher at higher redshifts; this is to be expected as, the further away the light source is, the further the light travels before reaching the observer, experiencing more lensing and thus producing a larger signal. It can also be seen that increasing noise reduces the accuracy, which is to be expected as the massive neutrino signal is distorted by the noise. The network is able to reach almost $100\%$ accuracy for a source redshift of 2.5 with $\ell_{\rm max}\simeq 5{,}000$. However, most interestingly, the accuracy of the network drops to almost $50\%$ after scaling out the 2-point information from the maps for all source redshifts and scales, except for low noise cases at small scales where it reaches $\sim 60\%$. This implies that the network is almost unable to tell the difference between the 0eV and 0.1eV maps, suggesting that most of the field-level information can be found using the power spectrum alone. One notable exception is in the case of $z=0.5$ where the loss of accuracy from scaling out the 2-point information is relatively small on small scales, implying that – even though the total information is relatively small (implied by the low accuracy) – a large fraction of this information exists beyond the 2-point.
@@
where $L(M_{\nu}|\kappa_{\rm obs})$ is the likelihood and $\kappa_{\rm obs}$ is an observed map from a cosmology with true neutrino mass corresponding to the ‘1 label’ in the classification; here $0.1\,{\rm eV}$ [64]. The lower row of Figure 3 shows how the difference in log likelihood $\Delta\ln L$ varies with $k_{\rm max}$, $z$, and $\sigma_{\kappa}$, for observations with $M_{\nu}=0.1\,{\rm eV}$ (taken from the test set). The overall trends are similar to the classification accuracy, with higher $k_{\rm max}$, higher $z$, and lower $\sigma_{\kappa}$, leading to a higher likelihood difference. In the most optimistic case of $z=2.5$ and $k_{\rm max}=2\pi\,h/{\rm Mpc}$ $(\ell_{\rm max}\simeq 10^{4})$ with low noise, the difference in log likelihood is $\Delta\ln L\simeq 4.5$. For a Gaussian likelihood this would correspond to $\Delta\chi^{2}\simeq 9$ and thus a $3\sigma$ detection.111 We note that, while it is instructive to consider the likelihood ratio to build intuition, this neglects any prior uncertainty on the neutrino mass as it only considers a $M_{\nu}=0.1{\rm eV}$ alternative hypothesis relative to a $M_{\nu}=0{\rm eV}$ null hypothesis. In a Bayesian analysis one would marginalize over a prior which spans a range of values of $M_{\nu}$, carefully taking into account volume effects [see e.g. 65, 66]. This in turn would reduce the significance of any detection of neutrino mass compared to the likelihood analysis here which corresponds to a delta-function prior at the observed value. This greatly decreases as lower source redshifts, lower scales, or higher noise is considered, dropping to $\Delta\ln L\simeq 0$. Most interestingly, it can also be seen that the likelihood difference is negligible when the 2-point information is removed in all cases, implying that there is little information beyond the 2-point.
@@
Figure 4 shows saliency maps using the Integrated Gradient method. This enables interpretation of what the network is learning and which parts of the map the model is most sensitive to. Each row displays a different weak lensing convergence map in the left column, followed by three saliency maps generated by the best three models from the optuna trials. Higher numbers in the saliency map implies that the network is paying more attention to that region of the map. We consider the case of $z=2.5$, $k_{\text{max}}={\pi}/{2}$ and $\sigma_{\kappa}=0.01$ where the accuracy of the network is high, and thus the saliency maps will be informative. It can be seen that there is much variability in the saliency maps, with the networks sometimes focusing more on underdense regions and sometimes on overdense regions. This implies that there is information in various features of the map, and that using an ensemble approach could allow us to extract more information, although this is beyond the scope of this work. To study the saliency more quantitatively, Figure 5 shows the average saliency in a pixel as a function of the (mean-subtracted and variance-scaled) convergence of the pixel $\tilde{\kappa}$. It can be seen that the saliency is highest for extreme values of $\tilde{\kappa}$, i.e. in the most overdense and underdense regions. This shows the complementarity of cluster-like and void-like regions, as found in 3-dimensional LSS examples which found complementarity between halo and void statistics [28, 29]. We find similar results in the noiseless case, but after increasing the noise to $\sigma_{\kappa}=0.1$ the higher saliency in the clusters is washed out. This is different to the trend observed for $\Omega_{m}$ and $\sigma_{8}$, where underdense regions were found to be more salient [54], showing promise for the prospect of breaking the degeneracy between these parameters.
@@ 4 Conclusions
This work quantified the ability of CNNs to extract information regarding neutrino mass from weak lensing convergence maps. We studied the information content as a function of scale cut, source redshift, and noise level for an LSST-like survey. We also investigated the effect of scaling out the 2-point information from the maps to study the beyond-2-point information. We found that a ResNet-18-based classifier was able to distinguish between a 0eV and 0.1eV neutrino mass cosmology with a high accuracy that improves as higher source redshifts and smaller scales are considered, reaching almost $100\%$ for a source redshift of 2.5 with $\ell_{\rm max}\simeq 5{,}000$. However, the accuracy of the network dropped to almost $50\%$ after scaling out the 2-point information from the maps, implying that most of the field-level information can be found using the power spectrum alone, in agreement with the simple cross-correlation arguments of [37]. We also went beyond using the neural network as a black box, by using Integrated Gradient saliency maps to understand which parts of the map the network is learning from. We found that the network pays most attention to the most overdense and underdense regions, with no clear preference for either. This suggests the two regions to be complementary, as found in studies that explicitly consider cluster-based and void-based summary statistics [28, 29].
@@
In this analysis we were restricted to using a classifier due to the availability of simulations, but in the future, when more simulations are produced, it would be interesting further work to do full parameter inference, studying the beyond 2-point information in a similar manner to [67] which considered $\Omega_{m}$ and $\sigma_{8}$ for halo clustering. While we have performed extensive hyperparameter tuning to facilitate the optimality of the CNN, it would also be interesting future work to study the field-level neutrino mass information using different architectures and methodologies, such as differentiable forward modelling to reconstruct the initial conditions of the Universe [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85] and normalizing flows [86, 87, 88]; although we do not expect this to change the conclusions. Furthermore, while we considered the information content in weak lensing convergence maps alone, it would be interesting to study the information when cross correlating convergence maps with other probes from LSS and the CMB using cross-survey simulations such as [89, 90] as this could provide a promising source of beyond-2-point information. Finally, it would be worthwhile to repeat this analysis while including baryonic effects to study the degeneracy between massive neutrinos and baryons at the field level [91].

### v3

@@ Abstract
Modern cosmological surveys probe the Universe deep into the nonlinear regime, where massive neutrinos suppress cosmic structure. Traditional cosmological analyses, which use the 2-point correlation function to extract information, are no longer optimal in the nonlinear regime, and there is thus much interest in extracting beyond-2-point information to improve constraints on neutrino mass. Quantifying and interpreting the beyond-2-point information is thus a pressing task. We study the field-level information in weak lensing convergence maps using convolution neural networks. We find that the network performance increases as higher source redshifts and smaller scales are considered – investigating up to a source redshift of 2.5 and $\ell_{\rm max}\simeq 10^{4}$ – verifying that massive neutrinos leave a distinct effect on weak lensing. However, the performance of the network significantly drops after scaling out the 2-point information from the maps, implying that most of the field-level information can be found in the 2-point correlation function alone. We quantify these findings in terms of the likelihood ratio and also use Integrated Gradient saliency maps to interpret which parts of the map the network is learning the most from. We find that, in the absence of noise, the network extracts a similar amount of information from the most overdense and underdense regions. However, upon adding noise, the information in underdense regions is distorted as noise disproportionately washes out void-like structures.
@@ 1 Introduction
One particularly promising method to measure neutrino mass is by studying weak gravitational lensing. This involves analyzing the distortions of light emitted from distant galaxies caused by the gravitational influence of LSS between the galaxy and the telescope. Traditionally, information is extracted from weak-lensing convergence maps using the power spectrum, or 2-point correlation function, which is optimal in the linear regime. However, upcoming surveys will probe progressively smaller, nonlinear, scales, in which regime the 2-point statistics are no longer guaranteed to be optimal. Various simulations have been developed to understand the effects of neutrinos on nonlinear scales [16, 17, 18, 19, 20, 21, 22], and have suggested that there is much information to be found on these nonlinear scales using higher-order statistics such as voids and wavelets. For example, when considering the 3-dimensional matter field it has been shown that higher-order statistics can improve neutrino mass constraints by a factor of $\sim 100$ compared to the power spectrum alone [23, 24, 25, 26, 27, 28, 29, 30], while for weak lensing more modest numbers are reported [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Understanding and interpreting the source of this information is thus a pressing task.
@@ 2.1 Data
    Cutting in Fourier Space: The images were transformed to Fourier space and cut at various scales in factors of 2 of the Nyquist, $k_{\rm max}=2\pi,\pi,\pi/2,\pi/4,\pi/8,\pi/16\,({h/{\rm Mpc})}$, and then transformed back into configuration space. As the maps are produced using the flat-sky approximation, we use a 2-dimensional wavenumber, denoted $k$, to represent the scale. To connect with curved-sky intuition, we show the relation between $k$ and multipole $\ell$ in Figure 2 for the various source redshifts considered.
@@
    Addition of Gaussian Noise: Gaussian noise with zero mean was added to mimic observational noise. We considered two standard deviations of $\sigma_{\kappa}=0.01$ and 0.1 (in the same units as $\kappa$), which is around the order of magnitude of noise expected for an LSST-like survey with a galaxy number density of $n_{\rm gal}\simeq 5\,{\rm arcmin}^{-2}$ (for one tomographic bin) and a shape noise of $\sigma_{\lambda}\simeq 0.3$ [32]. We also consider the noiseless maps.
@@ 2.2 Neural Network
We modified the ResNet-18 for our applications to take only 1 channel of input, instead of the default 3. We also replaced the final layer of the network by 2 fully connected layers—each followed by dropout and a LeakyReLU activation function—and then a third fully connected layer to output a scalar via a sigmoid activation function. We then apply a binary cross entropy loss function to train a classifier between the 0eV and 0.1eV neutrino mass simulations. For optimization we used the ADAM optimizer with an exponential learning rate scheduler, using early stopping to avoid overfitting. We report results for a batch size of 32, but also found that using a batch size of 64 or 128 made no improvement to the model. We use the pre-trained ResNet-18 weights as the starting point for training, but then refit all the weights during training. The data set was divided into training, validation, and testing sets according to a 70:15:15 split.
@@ 3 Results
Figure 3: Top: classification accuracy evaluated on the test set as a function of $k_{\text{max}}$. Bottom: difference in log likelihood between a cosmology with $M_{\nu}=0.1$ and $0\,{\rm eV}$ for observations with $M_{\nu}=0.1\,{\rm eV}$ (taken from the test set) as a function of $k_{\text{max}}$. Each panel and base color corresponds to a different source redshift $z$. The difference in color shades corresponds to the noise $\sigma_{\kappa}$, with lighter shades corresponding to higher noise. Solid lines represent the accuracy for a model trained on the full convergence maps, while dashed lines are for maps that have had the 2-point information scaled out. It can be seen for full maps that the accuracy and difference in likelihood are typically higher at smaller scales, higher redshifts, and lower noise. However, when the 2-point information is scaled out the accuracy and difference in likelihood are consistently significantly worse – almost $50\%$ and 0 respectively – indicating that much of the field-level information is contained in the 2-point alone.
@@
Figure 4: Visualization of the logarithm of the absoulte value of the Integrated Gradient saliency maps for the case of $z=2.5$, $k_{\text{max}}={\pi}/{2}$, and $\sigma_{\kappa}=0.01$ (without scaling out the 2-point information). Each row displays a different (mean-subtracted and variance-scaled) weak lensing convergence map in the left column, followed by three saliency maps generated by the best three models from the optuna trials. Higher numbers in the saliency map implies that the network is paying more attention to that region of the map. The saliency maps show a wide variability in terms of regions of the map they focus on, in some cases focusing on underdense regions, and in other cases focusing on overdense regions.
@@
Figure 5: Pixel-averaged values of the logarithm of the absolute value of the Integrated Gradient saliency maps as a function of the convergence map pixel value $\kappa$ (top row) and the standard-deviation-normalized convergence value $\tilde{\kappa}$ (bottom row), for models with $\sigma_{\kappa}=0$ (red) and $\sigma_{\kappa}=0.1$ (blue). The left column shows results computed using bin edges tailored to each noise level, while the right column compares both noise levels using a shared binning scheme derived from the $\sigma_{\kappa}=0$ case. All results correspond to models with $z=2.5$ and $k_{\text{max}}=\pi/2$, and include both massive and non-massive samples. The 2-point information has not been scaled out. Higher saliency values imply that the network is paying more attention to that region of the map, thus the network is paying most attention to extreme values of $\kappa$, i.e. the most overdense and underdense regions.
@@
The upper row of Figure 3 shows the classification accuracy evaluated on the test set for the different scale cuts $k_{\rm max}$, source redshifts $z$, and noise levels $\sigma_{\kappa}$, as well as the effect of scaling out the 2-point information. It can be seen that the accuracy of the network improves as smaller, nonlinear, scales are included; this is to be expected as the effect of massive neutrinos becomes more prominent on small scales. It can also be seen that the accuracy is higher at higher redshifts; this is to be expected as, the further away the light source is, the further the light travels before reaching the observer, experiencing more lensing and thus producing a larger signal. It can also be seen that increasing noise reduces the accuracy, which is to be expected as the massive neutrino signal is distorted by the noise. The network is able to reach almost $100\%$ accuracy for a source redshift of 2.5 with $\ell_{\rm max}\simeq 5{,}000$. However, most interestingly, the accuracy of the network drops to almost $50\%$ after scaling out the 2-point information from the maps for all source redshifts and scales, except for low noise cases at small scales where it reaches $\sim 60\%$. This implies that the network is almost unable to tell the difference between the 0eV and 0.1eV maps, suggesting that most of the field-level information can be found using the power spectrum alone. One notable exception is in the case of $z=0.5$ where the loss of accuracy from scaling out the 2-point information is relatively small on small scales, implying that – even though the total information is relatively small (implied by the low accuracy) – a large fraction of this information exists beyond the 2-point; this is especially the case in the zero-noise limit, implying that futuristic surveys with shape noise levels below LSST will have more information beyond the 2-point.
@@
where $L(M_{\nu}|\kappa_{\rm obs})$ is the likelihood and $\kappa_{\rm obs}$ is an observed map from a cosmology with true neutrino mass corresponding to the ‘1 label’ in the classification; here $0.1\,{\rm eV}$ [68]. The lower row of Figure 3 shows how the difference in log likelihood $\Delta\ln L$ varies with $k_{\rm max}$, $z$, and $\sigma_{\kappa}$, for observations with $M_{\nu}=0.1\,{\rm eV}$ (taken from the test set). The overall trends are similar to the classification accuracy, with higher $k_{\rm max}$, higher $z$, and lower $\sigma_{\kappa}$, leading to a higher likelihood difference. In the most optimistic case of $z=2.5$ and $k_{\rm max}=2\pi\,h/{\rm Mpc}$ $(\ell_{\rm max}\simeq 10^{4})$ with low noise, the difference in log likelihood is $\Delta\ln L\simeq 4.5$. For a Gaussian likelihood this would correspond to $\Delta\chi^{2}\simeq 9$ and thus a $3\sigma$ detection.111 We note that, while it is instructive to consider the likelihood ratio to build intuition, this neglects any prior uncertainty on the neutrino mass as it only considers a $M_{\nu}=0.1{\rm eV}$ alternative hypothesis relative to a $M_{\nu}=0{\rm eV}$ null hypothesis. In a Bayesian analysis one would marginalize over a prior which spans a range of values of $M_{\nu}$, carefully taking into account volume effects [see e.g. 69, 70]. This in turn would reduce the significance of any detection of neutrino mass compared to the likelihood analysis here which corresponds to a delta-function prior at the observed value. This greatly decreases as lower source redshifts, lower scales, or higher noise is considered, dropping to $\Delta\ln L\simeq 0$. Most interestingly, it can also be seen that the likelihood difference is negligible when the 2-point information is removed in all cases, implying that there is little information beyond the 2-point.
@@
Figure 4 shows saliency maps using the Integrated Gradient method. This enables interpretation of what the network is learning and which parts of the map the model is most sensitive to. Each row displays a different weak lensing convergence map in the left column, followed by three saliency maps generated by the best three models from the optuna trials. Higher numbers in the saliency map implies that the network is paying more attention to that region of the map. We consider the case of $z=2.5$ and $k_{\text{max}}={\pi}/{2}$, comparing models trained with $\sigma_{\kappa}=0$ and $\sigma_{\kappa}=0.1$. The corresponding saliency maps are expected to be informative. Looking at the Saliency plots for both $\sigma_{\kappa}=0$ and $\sigma_{\kappa}=0.1$ showed that the networks sometimes looked at both halo and void regions. This implies that there is information in various features of the map, and that using an ensemble approach could allow us to extract more information, although this is beyond the scope of this work.
To quantitatively analyze the saliency, Figure 5 shows the pixel-averaged logarithmic saliency as a function of the convergence value $\kappa$ (top) and normalized convergence $\tilde{\kappa}$ (bottom), when binned in terms of the $\kappa$ of each respective map (left) or with respect to the unnoised value $\kappa_{0}$ (right). In all cases, the saliency is largest in the most underdense and overdense regions of $\kappa$ or $\tilde{\kappa}$. For the noise-free case ($\sigma_{\kappa}=0$), there is slightly stronger saliency activation in halo regions, while for the noisy case ($\sigma_{\kappa}=0.1$), the activity appears more symmetrically distributed across the two tails. This shows the complementarity of cluster-like and void-like regions, as found in 3-dimensional LSS examples which found complementarity between halo and void statistics [28, 29]. When the saliency is binned in terms of the kappa value of the map in question (left column), the CNN pays equal attention to over and under-dense regions. However, when binning in terms of the unnoised convergence values (right column) the saliency of the noised map is severely washed out. This implies that in the presence of noise, the CNN pays the same attention to over and under-dense regions as it does when there is no noise, however, because many of the most underdense pixels are noise dominated, it is paying attention to noise, and thus information is lost. Underdense regions are less salient in the presence of noise as they have a lower $|\kappa_{0}|$ compared to overdense regions, and are thus washed out by the noise. This is similar to the trend observed for $\Omega_{m}$ and $\sigma_{8}$ by [58], where underdense regions were found to significantly lose information upon the addition of noise, however, our results are different to [58] as we find over and underdense regions are similarly salient in the absence of noise, while [58] found underdense regions to be more salient – implying a difference in what is a=most salient for $\Omega_{m},\sigma_{8},$ and $M_{\nu}$.
@@ 4 Conclusions
This work quantified the ability of CNNs to extract information regarding neutrino mass from weak lensing convergence maps. We studied the information content as a function of scale cut, source redshift, and noise level for an LSST-like survey. We also investigated the effect of scaling out the 2-point information from the maps to study the beyond-2-point information. We found that a ResNet-18-based classifier was able to distinguish between a 0eV and 0.1eV neutrino mass cosmology with a high accuracy that improves as higher source redshifts and smaller scales are considered, reaching almost $100\%$ for a source redshift of 2.5 with $\ell_{\rm max}\simeq 5{,}000$. However, the accuracy of the network dropped to almost $50\%$ after scaling out the 2-point information from the maps, implying that most of the field-level information can be found using the power spectrum alone, in agreement with the simple cross-correlation arguments of [41].
We also went beyond using the neural network as a black box, by using Integrated Gradient saliency maps to understand which parts of the map the network is learning from. We found that the network pays most attention to the most overdense and underdense regions. In the absence of noise, it pays similar attention to both overdense and underdense regions, with a slight preference for overdense regions. This complements findings from studies that explicitly considered cluster-based and void-based summary statistics and found them to be complementary [28, 29]. In the presence of noise, the network still pays attention to the most overdense and underdense pixels, however, many of these pixels are noise dominated, thus information is lost. This particularly affects underdense regions, which disproportionately lose information compared to overdense regions. This suggests that while both regions are complementary sources of information, underdense regions are more susceptible to information loss in noisy settings.
@@
In this analysis we were restricted to using a classifier due to the availability of simulations, but in the future, when more simulations are produced, it would be interesting further work to do full parameter inference, studying the beyond 2-point information in a similar manner to [71] which considered $\Omega_{m}$ and $\sigma_{8}$ for halo clustering. While we have performed extensive hyperparameter tuning to facilitate the optimality of the CNN, it would also be interesting future work to study the field-level neutrino mass information using different architectures and methodologies, such as differentiable forward modelling to reconstruct the initial conditions of the Universe [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89] and normalizing flows [90, 91, 92]; although we do not expect this to change the conclusions.
Furthermore, while we considered the information content in weak lensing convergence maps alone, it would be interesting to study the information when cross correlating convergence maps with other probes from LSS and the CMB using cross-survey simulations such as [93, 94], as this could provide a promising source of beyond-2-point information. It would also be fruitful to study the neutrino mass constraints while including systematic effects to investigate whether the relative constraining power of field-level to 2-point is modified. Finally, it would be worthwhile to repeat this analysis while including baryonic effects to study the degeneracy between massive neutrinos and baryons at the field level [95].